The admission fee at an amusement park is $1.25 for children and $4.80 for adults. On a certain day, 234 people entered the park, and the admission fees collected totaled 683 dollars. How many children and how many adults were admitted?

Answer by rothauserc(4718)

let c be the number of children and a be the number of adults
:
1) a +c = 234
:
2) 4.80a +1.25c = 683
:
solve equation 1 for a
:
a = 234 -c
:
substitute for a in equation 2
:
4.80(234 -c) +1.25c = 683
:
1123.2 -4.80c +1.25c = 683
:
-3.55c = -440.2
:
c = 124
:
a = 234 -124 = 110
:
********************************************************
There are 110 adults and 124 children
:
check answer by substituting for a and c in equation 2
:
4.80(110) +1.25(124) = 683
:
528 +155 = 683
:
683 = 683
:
answer checks
*********************************************************
: